OpenStudy (anonymous):

if f(x)=2x^2 -8x+10 find f^-1(x)

OpenStudy (anonymous):

Does that mean the inverse of f? In other words if f^-1 is the inverse of f, if y = f(x), then x = f^-1(y). You wont be able to solve this function for x in a way that will work for all y. This is because the function is a parabola, so can be two values of x which correspond to the same value of y. You can find an inverse that works for a limited range of y by using the quadratic formula to solve \[y = 2x^2 - 8x + 10\] for x in terms of y, and choosing one of the solutions.

OpenStudy (anonymous):

can u pls explain it better? i dnt relly get it

OpenStudy (anonymous):

Suppose we had a function g(x) \[y = g(x) = 2x\] We could solve this for x \[x = \frac{ g(x) }{ 2 } = \frac{ y }{ 2 }\] The above would be the inverse function of g, \[x = g^{-1}(y) = \frac{ y }{ 2 }\] But if we had a different function of x \[y = h(x) = x^2\] Then there is no inverse function because there are two x values for every y \[x = \pm \sqrt{y}\] so if we wrote\[h^{-1}(y)\] If wouldn't be clear if that should be equal to the positive or negative value of x. But if we only use h(x) with x greater than zero, than we have an inverse \[h^{-1} (y) = +\sqrt{y}\] There are errors in my first post, it should say "for all x" and "limited range of x", and not y. Hopefully this is clearer

OpenStudy (anonymous):

thanks for your help but what i need is the steps involved in the working